OpenStudy (anonymous):

HELP! MEDAL REWARDED!! What are the solutions to the following system of equations? y = x^2 + 12x + 30 8x − y = 10 a) (−4, −2) and (2, 5) b) (−2, −4) and (2, 5) c) (−2, −4) and (5, 2) d) No Real Solutions

OpenStudy (anonymous):

\[y = x^2 + 12x + 30\\ 8x − y = 10\] change to \[y = x^2 + 12x + 30\\ y=8x-10\] then solve \[x^2+12x+30=8x-10\]

OpenStudy (anonymous):

x^2 + 40 = -4x ?